OpenStudy (anonymous):

Let f(x) = 2x2 − 9x and g(x) = -5x2 + 4x. Find f(x) − g(x).

OpenStudy (anonymous):

\[f(x)-g(x)=2x^2-9x-(-5x^2+4x)\]

OpenStudy (anonymous):

\[=2x^2-9x+5x^2-4x=7x^2-13x\]

OpenStudy (anonymous):

So\[f(x)-g(x)=7x^2-13x\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np